# Functional differential equations driven by càdlàg rough paths

### v1

@@ 3.2. Controlled RDEs
    where the implicit multiplicative constant depends only on $p$, $\|f\|_{C^{3}_{b}}\vee\|\tilde{f}\|_{C^{3}_{b}}$, $\|\alpha,\alpha^{\prime}\|_{X,p}$, $\|\tilde{\alpha},\tilde{\alpha}^{\prime}\|_{X,p}$, $K$, and $\|\mathbf{X}\|_{p}$.
@@
Fix $(s,t)\in\Delta_{T}$ and let $(Y,Y^{\prime})\in\mathcal{V}^{p}_{X}$. We note that
$\|F(Y)\|_{p,[s,t]}=\|f((\alpha,Y))\|_{p,[s,t]}$
@@
$\quad\leq\|f\|_{C^{2}_{b}}(1+\|\alpha\|_{p,[s,t]})(1+\|Y\||_{p,[s,t]})$
and applying Lemma 3.2 to the enlarged controlled path $((\alpha,Y),(\alpha^{\prime},Y^{\prime}))$,
@@
The growth conditions thus hold with constant $C_{F}=\|f\|_{C^{2}_{b}}$ up to a multiplicative constant, which depends only on $p$, $\|\alpha\|_{p}$ and $\|\alpha,\alpha^{\prime}\|_{X,p}$.
@@
and similarly for each summand of the norm, so the Lipschitz conditions hold with constant $C_{F,K,X,X}$, which depends only on $p$, $\|f\|_{C^{3}_{b}}$, $K$, for $K>0$, $\|\alpha,\alpha^{\prime}\|_{X,p}$, and $\|X\|_{p}$. ∎
@@
(i) The existence and uniqueness of the solution follows immediately from Lemma 3.4 and Theorem 2.3. For the a priori estimate, note that $|F_{0}(\alpha,y)|\leq\|f\|_{C^{2}_{b}}$ and $C_{F}\lesssim_{p}\|f\|_{C^{2}_{b}}(1+\|\alpha\|_{p}+\|\alpha,\alpha^{\prime}\|_{X,p})^{2}$.
@@
the growth conditions hold with constant equal to $\|f-\tilde{f}\|_{C^{2}_{b}}$ up to a multiplicative constant which depends only on $p$, $\|\alpha\|_{p}$ and $\|\alpha,\alpha^{\prime}\|_{X,p}$. Further, it follows from the proofs of [FZ18, Lemma 3.1 and Lemma 3.5] that the growth conditions hold for
@@
with constant equal to $|\alpha_{0}-\tilde{\alpha}_{0}|+\|\alpha-\tilde{\alpha}\|_{p}+\|\alpha,\alpha^{\prime};\tilde{\alpha},\tilde{\alpha}^{\prime}\|_{X,p}$ up to a multiplicative constant which depends only on $p$, $\|f\|_{C^{3}_{b}}$, $\|\alpha\|_{p}$, $\|\tilde{\alpha}\|_{p}$, $\|\alpha,\alpha^{\prime}\|_{X,p}$, $\|\tilde{\alpha},\tilde{\alpha}^{\prime}\|_{X,p}$.
@@
satisfies Assumption 2.1 (i). Here, $C_{F-\tilde{F}}=\|f-\tilde{f}\|_{C^{2}_{b}}+|\alpha_{0}-\tilde{\alpha}_{0}|+\|\alpha-\tilde{\alpha}\|_{p}+\|\alpha,\alpha^{\prime};\tilde{\alpha},\tilde{\alpha}^{\prime}\|_{X,p}$ up to a multiplicative constant, which depends only on $p$, $\|f\|_{C^{3}_{b}}\vee\|\tilde{f}\|_{C^{3}_{b}}$, $\|\alpha\|_{p}$, $\|\tilde{\alpha}\|_{p}$, $\|\alpha,\alpha^{\prime}\|_{X,p}$, $\|\tilde{\alpha},\tilde{\alpha}^{\prime}\|_{X,p}$. ∎
@@ 3.3. RDEs with discrete time dependence
    where the implicit multiplicative constant depends only on $p$, $\ell$, $\|f\|_{C^{3}_{b}}$, $K$, and $\|\mathbf{X}\|_{p}$.
@@
with $(\bar{Y},\bar{Y}^{\prime})=((Y,Y,\ldots,Y),(Y^{\prime},Y^{\prime},\ldots,Y^{\prime}))\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k(\ell+1)})$. It follows analogously to Lemma 3.2 that the functional satisfies Assumption 2.1 with constants depending additionally on $\ell$, that is, $C_{F}=\|f\|_{C^{2}_{b}}$ up to a multiplicative constant, which depends only on $p$ and $\ell$, and $C_{F,K,X,X}$ depends only on $p$, $\ell$, $\|f\|_{C^{3}_{b}}$, $K$, for $K>0$, and $\|X\|_{p}$. Note that it is sufficient to check Assumption 2.1 (ii) for $X=\tilde{X}$. We can thus apply Theorem 2.3 to show that there exists a unique solution to the RDE (3.3) on the interval $[0,r_{1})$. We now aim to solve the RDE (3.3) iteratively on the subintervals $[r_{i},r_{i+1})$, $i=1,\ldots,\ell$, with $r_{\ell+1}=T$. Given the solution on $[r_{i-1},r_{i})$, with $r_{0}=0$, the value $Y_{r_{i}}$ is determined by the jump of $\mathbf{X}$ at time $r_{i}$. We therefore consider $(y_{i},y_{i}^{\prime})\in\mathcal{V}^{p}_{X}([r_{i},r_{i+1}];\mathbb{R}^{k})$, where
@@
for $i=\ell$. Analogously to Lemma 3.4, we can show that the functional satisfies Assumption 2.1 with constants depending additionally on $\ell$, that is $C_{F}=\|f\|_{C^{2}_{b}}$ up to a multiplicative constant, which depends only on $p$ and $\ell$, see the definition of $(\alpha_{i},\alpha_{i}^{\prime})\in\mathcal{V}^{p}_{X}([r_{i},r_{i+1}])$, and $C_{F,K,X,X}$ depends only on $p$, $\ell$, $\|f\|_{C^{3}_{b}}$, $K$, for $K>0$, and $\|X\|_{p}$. Note that it is again sufficient to check Assumption 2.1 (ii) for $X=\tilde{X}$. We can thus again apply Theorem 2.3 to show that there exists a unique solution to the RDE (3.3) on the interval $[r_{i},r_{i+1})$, that is
@@
where the implicit multiplicative constant depends only on $p$, $\ell$, $\|f\|_{C^{3}_{b}}$, $K$, and $\|\mathbf{X}\|_{p}$.
@@
where the implicit multiplicative depends only on $p$, $\|f\|_{C^{3}_{b}}$, $K$ and $\|\mathbf{X}\|_{p}$, see the definition of $(\alpha_{i},\alpha_{i}^{\prime}),(\tilde{\alpha}_{i},\tilde{\alpha}_{i}^{\prime})\in\mathcal{V}^{p}_{X}([r_{i},r_{i+1}];\mathbb{R}^{ik})$, $(y_{i},y_{i}^{\prime}),(\tilde{y}_{i},\tilde{y}_{i}^{\prime})\in\mathcal{V}^{p}_{X}([r_{i},r_{i+1}];\mathbb{R}^{k})$.
@@ 3.4. RDEs with constant delay
    where the implicit multiplicative constant depends only on $p$, $\ell$, $r_{1}$, $T$, $\|f\|_{C^{3}_{b}}$, $K$, $L$, and $\|\mathbf{X}\|_{p}$.
@@
(i) The existence and uniqueness of a solution follows by iteratively applying Corollary 3.3 (i) to intervals of the length $r_{1}$.
@@
$y_{i;t}=y_{t}+Y_{ir_{1}-}-y_{ir_{1}-}+F_{ir_{1}-}(Y)X_{ir_{1}-,ir_{1}}+F^{\prime}_{ir_{1}}(Y,Y^{\prime})\mathbb{X}_{ir_{1}-,ir_{1}},\quad t\in[ir_{1},(i+1)r_{1}],$
@@
where the implicit multiplicative constant depends only on $p$, $\ell$, $\|f\|_{C^{2}_{b}}$, $K$, $L$, and $\|\mathbf{X}\|_{p}$, see the definition of $(\alpha,\alpha^{\prime}),(\tilde{\alpha},\tilde{\alpha}^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k\ell})$.
@@ 3.5. RDEs with variable delay
$(F(Y),F^{\prime}(Y,Y^{\prime}))=(f((Y,\alpha)),\mathrm{D}((Y,\alpha))(Y^{\prime},\alpha^{\prime})),$
@@
    where the implicit multiplicative constant depends only on $p$, $\varepsilon$, $T$, $\eta$, $\|f\|_{C^{3}_{b}}$, $K$, $L$, and $\|\mathbf{X}\|_{p}$.
@@ 4.1. Delayed martingales as rough paths
The aim of this subsection is to construct a rough path above the stochastic process $X=(X_{t})_{t\in[0,T]}$, defined as
@@
in the spirit of stochastic Itô integration. Notice, although $Z$ is a martingale, the stochastic process $X$ is, in general, not a martingale anymore and, thus, we cannot directly apply stochastic Itô integration to define a rough path lift of $X$. As a first step, we derive the quadratic co-variation of the stochastic process $X$ in the next lemma.
@@
Proof.
@@ Appendix A Local estimates for rough integration
Let $\mathbf{X}\in\mathcal{D}^{p}([0,T];\mathbb{R}^{d})$ for $p\in(2,3)$ and $(Y,Y^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$. Suppose that the non-anticipative functional $(F,F^{\prime})\colon\mathcal{V}^{p}_{X}([0,T];\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{k}))\to\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$ satisfies Assumption 2.1 (i) with some constant $C_{F}$. Then, we have the local estimate
$\|R^{\int_{0}^{\cdot}F(Y)\mathrm{d}\mathbf{X}}\|_{\frac{p}{2},[s,t]}\lesssim_{p}C_{F}(1+\|Y,Y^{\prime}\|_{X,p,[s,t]})^{2}(1+\|X\|_{p,[s,t]})^{2}\|\mathbf{X}\|_{p,[s,t]}$
for all $(s,t)\in\Delta_{T}$.
@@
Let $(V,V^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$, and set $\Xi_{u,v}:=V_{u}X_{u,v}+V^{\prime}_{u}\mathbb{X}_{u,v}$ and $\delta\Xi_{u,r,v}:=\Xi_{u,v}-\Xi_{u,r}-\Xi_{r,v}$ for $s\leq u<r<v\leq t$. We note that
$\Big{|}R_{u,v}^{\int_{0}^{\cdot}V\mathrm{d}X}\Big{|}$ $\leq\Big{|}\int_{u}^{v}V_{r}\,\mathrm{d}\mathbf{X}_{r}-\Xi_{u,v}\Big{|}+|V^{\prime}_{u}||\mathbb{X}_{u,v}|$
@@
$\delta\Xi_{u,r,v}=R_{u,r}^{V}X_{r,v}+V^{\prime}_{u,r}\mathbb{X}_{r,v},$
@@
$|\delta\Xi_{u,r,v}|\leq\|R^{V}\|_{\frac{p}{2},[u,r]}\|X\|_{p,[r,v]}+\|V^{\prime}\|_{p,[u,r]}\|\mathbb{X}\|_{\frac{p}{2},[r,v]}.$
It then follows from the generalized sewing lemma, see [FZ18, Theorem 2.5], that
$\|R^{\int_{0}^{\cdot}V\mathrm{d}\mathbf{X}}\|_{\frac{p}{2},[s,t]}\lesssim_{p}\|V,V^{\prime}\|_{X,p,[s,t]}\|\mathbf{X}\|_{p,[s,t]}.$
For $(V,V^{\prime})=(F(Y),F^{\prime}(Y,Y^{\prime}))$, using Assumption 2.1 (i), we therefore obtain the estimate. ∎
@@
$\quad\lesssim C_{F,K,\mathbf{X},\tilde{\mathbf{X}}}(|Y_{s}-\tilde{Y}_{s}|+\|Y,Y^{\prime};\tilde{Y},\tilde{Y}^{\prime}\|_{X,\tilde{X},p,[s,t]}+\|X-\tilde{X}\|_{p,[s,t]})(\|\mathbf{X}\|_{p,[s,t]}\vee\|\tilde{\mathbf{X}}\|_{p,[s,t]})$
@@
for all $(s,t)\in\Delta_{T}$, if $\|Y,Y^{\prime}\|_{X,p,[s,t]}$, $\|\tilde{Y},\tilde{Y}^{\prime}\|_{\tilde{X},p,[s,t]}\leq K$, for some $K>0$, where the implicit multiplicative constant depends only on $p$, $\|\mathbf{X}\|_{p}$ and $\|\tilde{\mathbf{X}}\|_{p}$.
@@
$\|R^{\int_{0}^{\cdot}V\mathrm{d}\mathbf{X}}-R^{\int_{0}^{\cdot}\tilde{V}\mathrm{d}\tilde{\mathbf{X}}}\|_{\frac{p}{2},[s,t]}\lesssim\|V,V^{\prime};\tilde{V},\tilde{V}^{\prime}\|_{X,\tilde{X},p,[s,t]}\|\mathbf{X}\|_{p,[s,t]}+\|\tilde{V},\tilde{V}^{\prime}\|_{\tilde{X},p,[s,t]}\|\mathbf{X};\tilde{\mathbf{X}}\|_{p,[s,t]},$
where the implicit multiplicative constant depends only on $p$, $\|\mathbf{X}\|_{p}$ and $\|\tilde{\mathbf{X}}\|_{p}$. For $(V,V^{\prime})=(F(Y),F^{\prime}(Y,Y^{\prime}))$, $(\tilde{V},\tilde{V}^{\prime})=(F(\tilde{Y}),F^{\prime}(\tilde{Y},\tilde{Y}^{\prime}))$, using Assumption 2.1 (ii), we therefore obtain the estimate. ∎

### v2

@@ 3.2. Controlled RDEs
    where the implicit multiplicative constant depends on $p$, $\|f\|_{C^{3}_{b}}\vee\|\tilde{f}\|_{C^{3}_{b}}$, $\|\alpha,\alpha^{\prime}\|_{X,p}$, $\|\tilde{\alpha},\tilde{\alpha}^{\prime}\|_{X,p}$, $K$, and $\|\mathbf{X}\|_{p}$.
@@
Fix $(s,t)\in\Delta_{T}$ and let $(Y,Y^{\prime})\in\mathcal{V}^{p}_{X}$. It is clear that $|F_{t}(Y)|\leq\|f\|_{C^{2}_{b}}$, and we note that
$|F_{t-,t}(Y)|\leq\|F(Y)\|_{p,[s,t]}=\|f((\alpha,Y))\|_{p,[s,t]}$
@@
$\quad\leq\|f\|_{C^{2}_{b}}(1+\|\alpha\|_{p,[s,t]})(1+\|Y\||_{p,[s,t]}),$
and it holds that $\|Y\|_{p,[s,t]}\leq\|Y\|_{p,[s,t)}+|Y_{t-,t}|$ as well as $\|Y\|_{p,[s,t]}\leq C_{p}(1+(|Y^{\prime}_{s}|+\|Y^{\prime}\|_{p,[s,t]})\cdot\|X\|_{p,[s,t]}+\|R^{Y}\|_{\frac{p}{2},[s,t]})$. Applying Lemma 3.2 to the enlarged controlled path $((\alpha,Y),(\alpha^{\prime},Y^{\prime}))$,
@@
The growth conditions thus hold with constant $C_{F}=\|f\|_{C^{2}_{b}}$ up to a multiplicative constant, which depends on $p$, $\|\alpha\|_{p}$ and $\|\alpha,\alpha^{\prime}\|_{X,p}$.
@@
and similarly for each summand of the norm, so the Lipschitz conditions hold with constant $C_{F,K,X,X}$, which depends on $p$, $\|f\|_{C^{3}_{b}}$, $K$, for $K>0$, $\|\alpha,\alpha^{\prime}\|_{X,p}$, and $\|X\|_{p}$. ∎
@@
(i) The existence and uniqueness of the solution follows immediately from Lemma 3.4 and Theorem 2.4. For the a priori estimate, note that $C_{F}\lesssim_{p}\|f\|_{C^{2}_{b}}(1+\|\alpha\|_{p}+\|\alpha,\alpha^{\prime}\|_{X,p})^{2}$.
@@
the growth conditions hold with constant equal to $\|f-\tilde{f}\|_{C^{2}_{b}}$ up to a multiplicative constant which depends on $p$, $\|\alpha\|_{p}$ and $\|\alpha,\alpha^{\prime}\|_{X,p}$. Further, it follows from the proofs of [FZ18, Lemma 3.1 and Lemma 3.5] that the growth conditions hold for
@@
with constant equal to $|\alpha_{0}-\tilde{\alpha}_{0}|+\|\alpha-\tilde{\alpha}\|_{p}+\|\alpha,\alpha^{\prime};\tilde{\alpha},\tilde{\alpha}^{\prime}\|_{X,p}$ up to a multiplicative constant which depends on $p$, $\|f\|_{C^{3}_{b}}$, $\|\alpha\|_{p}$, $\|\tilde{\alpha}\|_{p}$, $\|\alpha,\alpha^{\prime}\|_{X,p}$, $\|\tilde{\alpha},\tilde{\alpha}^{\prime}\|_{X,p}$.
@@
satisfies Assumption 2.1 (i). Here, $C_{F-\tilde{F}}=\|f-\tilde{f}\|_{C^{2}_{b}}+|\alpha_{0}-\tilde{\alpha}_{0}|+\|\alpha-\tilde{\alpha}\|_{p}+\|\alpha,\alpha^{\prime};\tilde{\alpha},\tilde{\alpha}^{\prime}\|_{X,p}$ up to a multiplicative constant, which depends on $p$, $\|f\|_{C^{3}_{b}}\vee\|\tilde{f}\|_{C^{3}_{b}}$, $\|\alpha\|_{p}$, $\|\tilde{\alpha}\|_{p}$, $\|\alpha,\alpha^{\prime}\|_{X,p}$, $\|\tilde{\alpha},\tilde{\alpha}^{\prime}\|_{X,p}$. ∎
@@ 3.3. RDEs with discrete time dependence
    where the implicit multiplicative constant depends on $p$, $\ell$, $\|f\|_{C^{3}_{b}}$, $K$, and $\|\mathbf{X}\|_{p}$.
@@
with $(\bar{Y},\bar{Y}^{\prime})=((Y,Y,\ldots,Y),(Y^{\prime},Y^{\prime},\ldots,Y^{\prime}))\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k(\ell+1)})$. It follows analogously to Lemma 3.2 that the functional satisfies Assumption 2.1 with constants depending additionally on $\ell$, that is, $C_{F}=\|f\|_{C^{2}_{b}}$ up to a multiplicative constant, which depends on $p$ and $\ell$, and $C_{F,K,X,X}$ depends on $p$, $\ell$, $\|f\|_{C^{3}_{b}}$, $K$, for $K>0$, and $\|X\|_{p}$. Note that it is sufficient to check Assumption 2.1 (ii) for $X=\tilde{X}$. We can thus apply Theorem 2.4 to show that there exists a unique solution to the RDE (3.3) on the interval $[0,r_{1})$. We now aim to solve the RDE (3.3) iteratively on the subintervals $[r_{i},r_{i+1})$, $i=1,\ldots,\ell$, with $r_{\ell+1}=T$. Given the solution on $[r_{i-1},r_{i})$, with $r_{0}=0$, the value $Y_{r_{i}}$ is determined by the jump of $\mathbf{X}$ at time $r_{i}$. We therefore consider $(y_{i},y_{i}^{\prime})\in\mathcal{V}^{p}_{X}([r_{i},r_{i+1}];\mathbb{R}^{k})$, where
@@
for $i=\ell$. Analogously to Lemma 3.4, we can show that the functional satisfies Assumption 2.1 with constants depending additionally on $\ell$, that is $C_{F}=\|f\|_{C^{2}_{b}}$ up to a multiplicative constant, which depends on $p$ and $\ell$, see the definition of $(\alpha_{i},\alpha_{i}^{\prime})\in\mathcal{V}^{p}_{X}([r_{i},r_{i+1}])$, and $C_{F,K,X,X}$ depends on $p$, $\ell$, $\|f\|_{C^{3}_{b}}$, $K$, for $K>0$, and $\|X\|_{p}$. Note that it is again sufficient to check Assumption 2.1 (ii) for $X=\tilde{X}$. We can thus again apply Theorem 2.4 to show that there exists a unique solution to the RDE (3.3) on the interval $[r_{i},r_{i+1})$, that is
@@
where the implicit multiplicative constant depends on $p$, $\ell$, $\|f\|_{C^{3}_{b}}$, $K$, and $\|\mathbf{X}\|_{p}$.
@@
where the implicit multiplicative depends on $p$, $\|f\|_{C^{3}_{b}}$, $K$ and $\|\mathbf{X}\|_{p}$, see the definition of $(\alpha_{i},\alpha_{i}^{\prime}),(\tilde{\alpha}_{i},\tilde{\alpha}_{i}^{\prime})\in\mathcal{V}^{p}_{X}([r_{i},r_{i+1}];\mathbb{R}^{ik})$, $(y_{i},y_{i}^{\prime}),(\tilde{y}_{i},\tilde{y}_{i}^{\prime})\in\mathcal{V}^{p}_{X}([r_{i},r_{i+1}];\mathbb{R}^{k})$.
@@ 3.4. RDEs with constant delay
    where the implicit multiplicative constant depends on $p$, $\ell$, $r_{1}$, $T$, $\|f\|_{C^{3}_{b}}$, $K$, $L$, and $\|\mathbf{X}\|_{p}$.
@@
(i) The existence and uniqueness of the solution follows by iteratively applying Corollary 3.3 (i) to intervals of the length $r_{1}$.
@@
$y_{i;t}=y_{t}+Y_{ir_{1}-}-y_{ir_{1}-}+F_{ir_{1}-}(Y)X_{ir_{1}-,ir_{1}}+F^{\prime}_{ir_{1}-}(Y,Y^{\prime})\mathbb{X}_{ir_{1}-,ir_{1}},\quad t\in[ir_{1},(i+1)r_{1}],$
@@
where the implicit multiplicative constant depends on $p$, $\ell$, $\|f\|_{C^{2}_{b}}$, $K$, $L$, and $\|\mathbf{X}\|_{p}$, see the definition of $(\alpha,\alpha^{\prime}),(\tilde{\alpha},\tilde{\alpha}^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k\ell})$.
@@ 3.5. RDEs with variable delay
$(F(Y),F^{\prime}(Y,Y^{\prime}))=(f((Y,\alpha)),\mathrm{D}f((Y,\alpha))(Y^{\prime},\alpha^{\prime})),$
@@
    where the implicit multiplicative constant depends on $p$, $\varepsilon$, $T$, $\eta$, $\|f\|_{C^{3}_{b}}$, $K$, $L$, and $\|\mathbf{X}\|_{p}$.
@@ 4.1. Delayed martingales as rough paths
The aim of this subsection is to construct a random rough path above the stochastic process $X=(X_{t})_{t\in[0,T]}$, defined as
@@
in the spirit of stochastic Itô integration. Notice, although $Z$ is a martingale, the stochastic process $X$ is, in general, not a martingale anymore and, thus, we cannot directly apply stochastic Itô integration to define a rough path lift of $X$. Moreover, if $Z$ is a semimartingale, it seems not obvious under which conditions the associated process $X$ is again a semimartingale.
As a first step to construct a random rough path above the stochastic process $X$, in the next lemma, we derive the existence of an auxiliary process, inspired by the quadratic co-variation of martingales.
@@
Remark 4.3.
The stochastic integral $\int_{0}^{t}Z^{i_{1}}_{u-r_{j_{1}}-}\,\mathrm{d}Z^{i_{2}}_{u-r_{j_{2}}}$ can be defined using classical stochastic Itô integration if $r_{j_{1}}>r_{j_{2}}$. Indeed, the stochastic process $(Z^{i_{2}}_{t-r_{j_{2}}})_{t\in[0,T]}$ is a martingale and the stochastic process $(Z^{i_{1}}_{t-r_{j_{1}}})_{t\in[0,T]}$ is predictable, both with respect to the filtration $(\mathcal{F}_{t-r_{j_{2}}})_{t\in[0,T]}$ with $\mathcal{F}_{t}:=\{\Omega,\emptyset\}$ for $t<0$.
Further, for $i\geq j$, we have that $\mathbb{X}^{ij}_{s,t}:=\int_{0}^{t}X^{i}_{u-}\,\mathrm{d}X^{j}_{u}-\int_{0}^{s}X^{i}_{u-}\,\mathrm{d}X^{j}_{u}-X^{i}_{s}X^{j}_{s,t}=\int_{s}^{t}X^{i}_{s,u-}\,\mathrm{d}X^{j}_{u}\\ =\int_{s}^{t}\int_{s}^{u-}\mathrm{d}X^{i}_{r}\,\mathrm{d}X^{j}_{u}$, that is, $\mathbb{X}^{ij}_{s,t}$ coincides with the $2$-fold iterated integral, for $(s,t)\in\Delta_{T}$.
Proof of Proposition 4.2.
@@ Appendix A Local estimates for rough integration
Let $\mathbf{X}\in\mathcal{D}^{p}([0,T];\mathbb{R}^{d})$ for $p\in(2,3)$ and $(Y,Y^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$. Suppose that the non-anticipative functional $(F,F^{\prime})\colon\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})\to\mathcal{V}^{p}_{X}([0,T];\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{k}))$ satisfies Assumption 2.1 (i) with some constant $C_{F}$. Then, we have the local estimate
$\|R^{\int_{0}^{\cdot}F(Y)\mathrm{d}\mathbf{X}}\|_{\frac{p}{2},[s,t]}\lesssim C_{F}(1+\|Y,Y^{\prime}\|_{X,p,[s,t]})^{2}(1+\|X\|_{p,[s,t]})^{2}\|\mathbf{X}\|_{p,[s,t]},$
for all $(s,t)\in\Delta_{T}$, where the implicit multiplicative constant depends only on $p$.
@@
Let $(V,V^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{k}))$, and set $\Xi_{u,v}:=V_{u}X_{u,v}+V^{\prime}_{u}\mathbb{X}_{u,v}$ and $\delta\Xi_{u,r,v}:=\Xi_{u,v}-\Xi_{u,r}-\Xi_{r,v}$ for $s\leq u<r<v\leq t$. Here, strictly speaking, in writing $V^{\prime}_{u}\mathbb{X}_{u,v}$, we use the canonical identification of $\mathcal{L}(\mathbb{R}^{d};\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{k}))$ with $\mathcal{L}(\mathbb{R}^{d}\otimes\mathbb{R}^{d};\mathbb{R}^{k})$. We note that
$\Big{|}R_{u,v}^{\int_{0}^{\cdot}V\mathrm{d}\mathbf{X}}\Big{|}$ $\leq\Big{|}\int_{u}^{v}V_{r}\,\mathrm{d}\mathbf{X}_{r}-\Xi_{u,v}\Big{|}+|V^{\prime}_{u}||\mathbb{X}_{u,v}|$
@@
$-\delta\Xi_{u,r,v}=R_{u,r}^{V}X_{r,v}+V^{\prime}_{u,r}\mathbb{X}_{r,v},$
@@
$|\delta\Xi_{u,r,v}|$
$\quad\leq\|R^{V}\|_{\frac{p}{2},[u,r]}\|X\|_{p,[r,v]}+\|V^{\prime}\|_{p,[u,r]}\|\mathbb{X}\|_{\frac{p}{2},[r,v]}$
$\quad=w_{1,1}(u,r)^{\frac{2}{p}}w_{2,1}(r,v)^{\frac{1}{p}}+w_{1,2}(u,r)^{\frac{1}{p}}w_{2,2}(r,v)^{\frac{2}{p}},$
where $w_{1,1}(s,t):=\|R^{V}\|_{\frac{p}{2},[s,t]}^{\frac{p}{2}}$, $w_{2,1}(s,t):=\|X\|_{p,[s,t]}^{p}$, $w_{1,2}(s,t):=\|V^{\prime}\|_{p,[s,t]}^{p}$, $w_{2,2}(s,t):=\|\mathbb{X}\|_{\frac{p}{2},[s,t]}^{\frac{p}{2}}$, $(s,t)\in\Delta_{T}$, are control functions and $\frac{1}{p}+\frac{2}{p}>1$. It then follows from the generalized sewing lemma, see [FZ18, Theorem 2.5], that
$\|R^{\int_{0}^{\cdot}V\mathrm{d}\mathbf{X}}\|_{\frac{p}{2},[s,t]}$
$\quad\lesssim(\|R^{V}\|_{\frac{p}{2},[s,t]}\|X\|_{p,[s,t]}+\|V^{\prime}\|_{p,[s,t]}\|\mathbb{X}\|_{\frac{p}{2},[s,t]}+(|V^{\prime}_{s}|+\|V^{\prime}\|_{p,[s,t]})\|\mathbb{X}\|_{\frac{p}{2},[s,t]})$
$\quad\lesssim\|V,V^{\prime}\|_{X,p,[s,t]}\|\mathbf{X}\|_{p,[s,t]},$
where the implicit multiplicative constant depends only on $p$. For $(V,V^{\prime})=(F(Y),F^{\prime}(Y,Y^{\prime}))$, using Assumption 2.1 (i), we therefore obtain the estimate. ∎
@@
$\quad\lesssim C_{F,K,X,\tilde{X}}(|Y_{s}-\tilde{Y}_{s}|+\|Y,Y^{\prime};\tilde{Y},\tilde{Y}^{\prime}\|_{X,\tilde{X},p,[s,t]}+\|X-\tilde{X}\|_{p,[s,t]})(\|\mathbf{X}\|_{p,[s,t]}\vee\|\tilde{\mathbf{X}}\|_{p,[s,t]})$
@@
for all $(s,t)\in\Delta_{T}$, if $\|Y,Y^{\prime}\|_{X,p,[s,t]}$, $\|\tilde{Y},\tilde{Y}^{\prime}\|_{\tilde{X},p,[s,t]}\leq K$, for some $K>0$, where the implicit multiplicative constant depends on $p$, $\|\mathbf{X}\|_{p}$ and $\|\tilde{\mathbf{X}}\|_{p}$.
@@
$\|R^{\int_{0}^{\cdot}V\mathrm{d}\mathbf{X}}-R^{\int_{0}^{\cdot}\tilde{V}\mathrm{d}\tilde{\mathbf{X}}}\|_{\frac{p}{2},[s,t]}$
$\quad\lesssim_{p}(1+\|\mathbf{X}\|_{p,[s,t]}+\|\tilde{\mathbf{X}}\|_{p,[s,t]})(\|V,V^{\prime};\tilde{V},\tilde{V}^{\prime}\|_{X,\tilde{X},p,[s,t]}\|\mathbf{X}\|_{p,[s,t]}$
$\qquad+\|\tilde{V},\tilde{V}^{\prime}\|_{\tilde{X},p,[s,t]}\|\mathbf{X};\tilde{\mathbf{X}}\|_{p,[s,t]})$
$\quad\lesssim\|V,V^{\prime};\tilde{V},\tilde{V}^{\prime}\|_{X,\tilde{X},p,[s,t]}\|\mathbf{X}\|_{p,[s,t]}+\|\tilde{V},\tilde{V}^{\prime}\|_{\tilde{X},p,[s,t]}\|\mathbf{X};\tilde{\mathbf{X}}\|_{p,[s,t]},$
where the implicit multiplicative constant depends on $p$, $\|\mathbf{X}\|_{p}$ and $\|\tilde{\mathbf{X}}\|_{p}$. For $(V,V^{\prime})=(F(Y),F^{\prime}(Y,Y^{\prime}))$, $(\tilde{V},\tilde{V}^{\prime})=(F(\tilde{Y}),F^{\prime}(\tilde{Y},\tilde{Y}^{\prime}))$, using Assumption 2.1 (ii), we therefore obtain the estimate. ∎